OpenStudy (anonymous):

solve the given differential equation by using appropriate solutions x*dx+(y-2x)dy=0

OpenStudy (anonymous):

are bernoulli's homogenious

OpenStudy (experimentx):

wasn't Bernoulli's exactly ... it was first order linear differential equation.

OpenStudy (anonymous):

I cant view the whole link

OpenStudy (anonymous):

I know I have to use substitution, but I don't know what to substitute, the answer is attatched below

OpenStudy (experimentx):

my previous answer was also incorrect, should have seen the equation more clearly,

OpenStudy (experimentx):

change y=vx, dy/dx = xdv/dx + v

OpenStudy (experimentx):

also cancel out, x's on the ratio, ... then use variable separation method.

OpenStudy (anonymous):

i dont get it could you walk me through it

OpenStudy (experimentx):

dy/dx+x/(y-2x)=0 put y = vx then you will have dy/dx = xdv/dx + v or xdv/dx + v +x/(vx-2x) = 0 now solve using variable separation method.

OpenStudy (anonymous):

wouldnt it be (y-2x)/x ?

OpenStudy (experimentx):

?? y is changed

OpenStudy (experimentx):

in to vx => cancel out that x's on second term

OpenStudy (anonymous):

I assume I can integrate now?

OpenStudy (experimentx):

after separating variables, v and x

OpenStudy (anonymous):

this seems good?

OpenStudy (anonymous):

before the integration

OpenStudy (experimentx):

yeah .. must be something like that.

OpenStudy (anonymous):

im not getting teh answer

OpenStudy (anonymous):

that they have in book

OpenStudy (experimentx):

change v back to y/x

OpenStudy (anonymous):

this is a homogeneous differential equation, you just need to substitute\[y = zx\]and then the it will become a separable equation

OpenStudy (anonymous):

can some one solve this for me because I have abosolutely no clue

OpenStudy (anonymous):

is it x dx or x*dx???

OpenStudy (anonymous):

x dx

OpenStudy (anonymous):

ok let me if i can help u

OpenStudy (anonymous):

\[xdx+\left( y-2x \right)dy=0\]\[\frac{dy}{dx}=\frac{-x}{y-2x}\]let\[y=zx\]\[\frac{dy}{dx}=\frac{dz}{dx}x + z\]and then substitute to the equation\[\frac{dz}{dx}x+z=\frac{-x}{zx - 2x}\]\[\frac{dz}{dx}x=\frac{-z^{2}+2z-1}{z+2}\]\[\frac{z+2}{-z^{2} + 2z - 1}dz=\frac{dx}{x}\]now you just need to integrate both sides